# TrustGeoGen: Scalable and Formal-Verified Data Engine for Trustworthy Multi-modal Geometric Problem Solving

### v1

@@ 3.2 Deep Reasoning Augmentation
As mentioned earlier, the randomly constructed scenarios by TrustGeoGen exhibit a rapid decline in reasoning length within the ultra-long range, posing challenges for generating deep-reasoning problems. To address this, we selected 226 samples with the longest reasoning lengths from an initial dataset of 200K and applied bootstrap augmentation to increase the proportion of deep-reasoning problems. Finally we obtained 376 geometric scenes after applying the bootstrap method. As shown in Fig. 3(c), after the bootstrap process, the data distribution in the region where reasoning length is smaller than 40 decreased, while it increased in the region where reasoning length is larger than 40. This method can be repeatedly applied to efficiently construct reasoning problems with significant depth and complexity.
@@ 3.3 Construction and Analysis of GeoTrust-Test
The synthetic data generated by TrustGeoGen can also be utilized to evaluate a model’s capabilities in the domain of deep geometric reasoning. From an initial dataset of 200K samples, we manually curated a set of 240 high-quality problems as the test set, graded by different levels of difficulty. As shown in Fig. 4, these problems are divided into four tiers, with each tier containing 60 problems. The reasoning lengths for $Tier_{1}$ range from 5 to 10 steps, $Tier_{2}$ spans from 10 to 20 steps, $Tier_{3}$ ranges from 20 to 50 steps, and $Tier_{4}$ exceeds 50 steps. It is worth noting that, to introduce distractors into the questions, the premises provided may not necessarily all be utilized in the reasoning process, resulting in a potentially lower premise ratio.
4 Experiment
@@ 4.1 Dataset, Metric, and Implementation Detail
Dataset. To evaluate the capability of MLLMs in solving complex geometry problems, we construct GeoTrust-test by partitioning the original GeoTrust dataset. As introduced in Sec. 3.3, the testset comprehensively covers four difficulty levels ranging from $Tier_{1}$ to $Tier_{4}$. To validate the effectiveness of our data construction, we additionally sample 3K geometry problems as GeoTrust-train, with each difficulty tier ($Tier_{1}$ to $Tier_{3}$§§§Considering the maximum output token limit of the trained open-source models, we excluded data from $Tier_{4}$.) containing 1K instances. GeoQA [6] is employed to evaluate the OOD generalization performance, comprising 4,998 geometric problems and solutions in natural language.
@@ 4.2 Performance of Existing MLLMs
As shown in Tab. 1, the experimental results from the GeoTrust-test testset reveal critical insights into the capabilities of MLLMs in tackling complex geometric problems. Existing open-source models, such as Qwen2.5-VL-7B-Instruct [5], demonstrate significant limitations, achieving only a 12.92% overall accuracy (31/240), which also underscores the inherent challenges of the dataset and confirms its independence from prior biases in existing open-source testsets. Besides, closed-source commercial models like OpenAI-o1 [25] exhibit stronger reasoning capabilities, solving 109 problems (45.42% accuracy), though their performance sharply declines on higher-difficulty $Tier_{4}$ questions (18.33% accuracy), highlighting persistent gaps in handling advanced geometric reasoning. Notably, the open-source reasoning model DeepSeek-R1For DeepSeek-R1, we only provide natural language questions. [10] matches OpenAI-o1’s overall accuracy, bridging the performance gap between open and closed-source systems and emphasizing the potential of scalable architectures. A consistent trend across all models is the progressive degradation of accuracy as problem difficulty escalates, validating the rigorous tiered difficulty framework of GeoTrust-test. These findings collectively emphasize the need for enhanced reasoning architectures and training paradigms to address the steep complexity gradient in geometric problem solving.
4.3 Advantages Provided by Trustworthy Data
Table 2: Comparison of performance on GeoTrust-test using trustworthy GeoTrust-train data and pseudo-labels annotated by OpenAI-o1 [25]. All trained on Qwen2-VL-7B-Instruct [5].
Training Data Total $Tier_{1}$ $Tier_{2}$ $Tier_{3}$
- 7 5 2 0
pseudo-labels 21 (+14) 13 (+8) 3 (+1) 5 (+5)
100% GeoTrust-train 38 (+31) 27 (+22) 6 (+4) 5 (+5)
50% GeoTrust-train 27 (+20) 22 (+17) 4 (+2) 1 (+1)
30% GeoTrust-train 12 (+5) 10 (+5) 2 (-) 0 (-)
To validate the effectiveness of the trustworthy data generated by TrustGeoGen, we compared our data with the ones generated by OpenAI-o1 [25]. We collect pesudo-labels generated by the OpenAI-o1 [25] via prompting the model with the problems in GeoTrust-train. We train Qwen2-VL-7B-Instruct [36] on both GeoTrust-train (with 100%, 50% and 30% data) and pesudo-labels, and report the number of problems solved from $Tier_{1}$ to $Tier_{3}$. As shown in  Tab. 2, the model trained on 100% of GeoTrust-train outperforms the model trained on pesudo-labels, which demonstrates that trustworthy data hold substantial advantages over SOTA-generated pseudo-labels under equivalent data amount. Besides, as the amount of GeoTrust-train data increases, the model’s overall performance continues to improve, which further validates the effectiveness of our dataset. However, while the improvements on $Tier_{1}$ are notable, the gains on $Tier_{2}$ and $Tier_{3}$ remain significantly less pronounced, suggesting that supervised fine-tuning (SFT) does not produce substantial effects on more challenging geometric reasoning tasks.
@@ 4.4 OOD Generalization
The experiments are conducted on three pre-trained vision-language models: LLaVA-1.5-7B [19], LLaVA-1.5-13B [19] and Qwen2-VL-7B-Instruct [36]. For each model, we evaluate on GeoQA [6] with two training settings: (1) fine-tuning solely on the GeoQA dataset, and (2) fine-tuning with a combination of GeoQA and GeoTrust-train datasets to explore generalization. The results in Tab. 3 demonstrate that incorporating GeoTrust-train data not only achieves substantial improvements over the original pre-trained models (e.g., +11.81% for LLaVA-1.5-7B and +5.57% for Qwen2-VL-7B ), but also surpasses models fine-tuned solely on the in-distribution GeoQA training set. For example, LLaVA-1.5-7B improves from 25.73% accuracy (GeoQA-only fine-tuning) to 28.12% when augmented with GeoTrust-train, marking a +2.39% absolute gain over the in-distribution baseline, highlighting the generalization capability of our data.
Table 3: OOD Generalization on GeoQA [6] when GeoTrust-train is leveraged for training, where $\Delta$ indicates the improvement compared to the pre-trained baseline.
Model Training Data acc $\Delta$
LLaVA-1.5-7B [19] - 16.31% -
GeoQA 25.73% +9.42%
GeoQA+GeoTrust-train 28.12% +11.81%
LLaVA-1.5-13B [19] - 18.83% -
GeoQA 29.84% +11.01%
GeoQA+GeoTrust-train 33.16% +14.33%
Qwen2-VL-7B-Instruct [36] - 37.67% -
GeoQA 41.11% +3.44%
GeoQA+GeoTrust-train 43.24% +5.57%

### v2

@@ 4.2 Deep Reasoning Augmentation
As mentioned earlier, the randomly constructed scenarios by TrustGeoGen exhibit a rapid decline in reasoning length within the ultra-long range, posing challenges for generating deep-reasoning problems. To address this, we selected 226 samples with the longest reasoning lengths from an initial dataset of 200K and applied bootstrap augmentation to increase the proportion of deep-reasoning problems. Finally we obtained 376 geometric scenes after applying the bootstrap method. As shown in  fig.˜4(c), after the bootstrap process, the data distribution in the region where reasoning length is smaller than 40 decreased, while it increased in the region where reasoning length is larger than 40. This method can be repeatedly applied to efficiently construct reasoning problems with significant depth and complexity.
@@ 4.3 Construction and Analysis of GeoTrust-Test
The synthetic data generated by TrustGeoGen can also be utilized to evaluate a model’s capabilities in the domain of deep geometric reasoning. From an initial dataset of 200K samples, we manually curated a set of 240 high-quality problems as the test set, graded by different levels of difficulty. Unlike the training data, which incorporates a mix of proof and solution-type problems, the GeoTrus-test is composed exclusively of problems with numerical solutions to simplify the evaluation process. As shown in  fig.˜5, these problems are divided into four tiers, with each tier containing 60 problems. The reasoning lengths for $Tier_{1}$ range from 5 to 10 steps, $Tier_{2}$ spans from 10 to 20 steps, $Tier_{3}$ ranges from 20 to 50 steps, and $Tier_{4}$ exceeds 50 steps. It is worth noting that, to introduce distractors into the questions, the premises provided may not necessarily all be utilized in the reasoning process, resulting in a potentially lower premise ratio.
5 Experiments
@@
  • Whether the data generated by TrustGeoGen have sufficient complexity to provide gains for SOTA models?
@@
TABLE III: Complexity comparison between different benchmarks. The most difficult results are highlighted in bold, and the second difficult results are highlighted with an underline. It is indicated that our geometry constructed problems are more challenging for most mainstream MLLMs.
Model Release Date
GeoQA
(mid-school level)
Geometry3K
(mid-school level)
OlympiadBench-Geo
(olympiad level)
GeoTrust-test
GPT-4o openai2024gpt4o May, 2024 42.31% 31.45% 13.39% 25.83%
Claude-3.7-sonnet (2024claude, ) Feb, 2025 49.73% 33.28% 17.86% 27.50%
Qwen2.5-VL-72B Qwen2.5-VL Jan, 2025 67.90% 35.44% 29.46% 28.33%
Intern-S1 bai2025intern Jul, 2025 62.47% 52.31% 49.11% 43.33%
Gemini-2.5-procomanici2025gemini Jun, 2025 79.58% 80.70% 75.00% 43.33%
OpenAI-o3 openai_o3 Apr, 2025 79.31% 81.03% 77.68% 45.83%
@@ 5.1 Dataset, Metric, and Implementation Detail
Dataset. To evaluate the capability of MLLMs in solving complex geometry problems, we construct GeoTrust-test by partitioning the original GeoTrust dataset. As introduced in section˜4.3, the testset GeoTrust-test comprehensively covers four difficulty levels ranging from $Tier_{1}$ to $Tier_{4}$. To validate the effectiveness of our data construction, we additionally sample 2342 geometry problems as GeoTrust-train, with different difficulty tier.
@@ 5.2 Performance of Existing MLLMs
As shown in table˜II, the experimental results from the GeoTrust-test testset reveal critical insights into the capabilities of MLLMs in tackling complex geometric problems. Existing open-source models, such as Qwen2-VL-7B, demonstrate significant limitations, achieving only a 4.58% overall accuracy (11/240), which also underscores the inherent challenges of the dataset and confirms its independence from prior biases in existing open-source testsets. The closed-source reasoning model, OpenAI-o3, demonstrated the strongest geometric reasoning capability, solving 110 problems with an accuracy of 45.83%. Furthermore, models that have undergone "deep thinking" training (e.g., OpenAI-o3, Gemini-2.5-pro, Intern-S1, DeepSeek-R1) exhibited a more gradual decline in accuracy as problem difficulty increased (from Tier1 to Tier4). In contrast, models without such specialized training (e.g., GPT-4o, Claude-3.7-sonnet, Qwen2.5-VL-72B) experienced a sharp drop in accuracy on more challenging problems. This distinction underscores the critical importance of deep reasoning abilities for GPS. The consistent decrease in accuracy across all models with rising difficulty collectively emphasizes the need for enhanced reasoning architectures and training paradigms to address the steep complexity gradient in geometric problem solving.
5.3 Complexity Analysis Between Benchmarks
Although constructing geometric data using formal engines can enhance acquisition efficiency, it presents challenges in ascertaining the data’s complexity and its actual contribution to model training. To address this, TrustGeoGen utilizes filtering parameters, $\tau_{l}$ and $\tau_{r}$, to screen for problems of varying difficulty, thereby enabling the selective generation of data that provides tangible benefits to the model. As detailed in table˜III, we evaluated several state-of-the-art (SOTA) multimodal large language models on four distinct geometry benchmarks to assess their respective complexity. The performance of SOTA models indicates that GeoTrust-test presents a level of difficulty comparable to OlympiadBench-Geo, with models generally struggling to achieve high scores on either. Notably, even well-pretrained models like Gemini-2.5-pro and OpenAI-o3, which exhibit substantially improved accuracy on OlympiadBench-Geo thanks to extensive pre-training on geometry data, still fail to surpass the 50% accuracy mark on GeoTrust-test. This underscores the potential of TrustGeoGen to further boost the performance of these advanced models by providing a continuous stream of fresh and challenging data.
5.4 Effectiveness Analysis of Trustworthy Data
TrustGeoGen constructs trustworthy geometric reasoning data using a formal engine. It bridges the gap between formal and human-like reasoning through "connection thinking" (algorithm˜1). Furthermore, thinking template data—comprising multi-solution data (section˜3.3) and self-reflective traceback data (section˜3.4)—is utilized to introduce diverse reasoning pathways. In this section, we design comprehensive experiments to validate the effectiveness of our trustworthy data.
TABLE IV: Translation performance comparison across different models and training data sources. The $\Delta$ column indicates the improvement over the "Pretraining" baseline. Improvements are marked with an upward arrow ($\uparrow$).
Training Data LLaVA-1.5-7B LLaVA-1.5-13B Qwen2-VL-2B Qwen2-VL-7B
Accuracy $\Delta$ Accuracy $\Delta$ Accuracy $\Delta$ Accuracy $\Delta$
Pretraining 1.67% (4) - 2.08% (5) - 3.33% (8) - 4.58% (11) -
[…]
5.4.1 Necessary of Connection Thinking
We conducted experiments to validate our data’s effectiveness using four models: LLaVA-1.5-7B, LLaVA-1.5-13B, Qwen2-VL-2B, and Qwen2-VL-7B. Each model was fine-tuned via Supervised Fine-Tuning (SFT) on a sampled dataset of 2,158 instances. We applied two distinct training strategies separately: one using the NL translation data and the other using the connection thinking data, both of which are detailed in fig.˜2. All fine-tuned models were subsequently evaluated on the GeoTrust-test benchmark.
As shown in table˜IV, the results demonstrate that both training approaches lead to a significant increase in model accuracy. Notably, the models trained with the connection thinking data exhibited a more substantial improvement, consistently achieving accuracy gains of over 10%.
Furthermore, we analyzed the outputs of models trained on different datasets. As illustrated in fig.˜6, there is a substantial difference in the content generated by models trained on the two respective datasets. The model trained on the "NL Translation" data tends to produce repetitive content and exhibits apparent logical fallacies. In contrast, the model trained on our proposed "connection thinking" data generates outputs that are significantly more coherent and logically sound.
The "NL Translation" data suffers from two key limitations that prevent a model from truly acquiring reasoning capabilities. First, its formal-language reasoning steps are highly structured, merely presenting premises, applied theorems, and derived conclusions in a templated format without elucidating the underlying rationale of the proof process. A model trained on such data learns to replicate theorems it has seen rather than how to apply them proficiently. Second, a significant gap exists between this formal reasoning process and authentic human-like reasoning, as there are no explicit logical connections between adjacent steps. Consequently, each subsequent step appears to emerge abruptly, lacking logical coherence.
To address these limitations, our "connection thinking" approach uses the "NL Translation" data as a foundational skeleton but augments it by inserting a detailed thought process between each pair of consecutive reasoning steps. This enrichment not only explains each step in detail but also articulates the logical link between them by considering previously established conclusions and the ultimate goal. This enriched data structure enables the model to master the application of plane geometry theorems, thereby significantly improving its geometric problem-solving capabilities.
Figure 6: Fragments of model’s response: The response of model trained on NL Translation dataset is repeating & illogical, while the response of model trained on Connection Thinking dataset is reasonable and clear.
5.4.2 Advantages Provided by Trustworthy Data
To validate the effectiveness of data generated by TrustGeoGen, we conducted a comparative experiment using pseudo-labeled data. We fed 2,158 questions generated by TrustGeoGen into OpenAI-o3 openai_o3 and obtained answers from the model. The responses were then evaluated using the metrics described in section˜5.1, resulting in 846 correctly answered questions, which we treat as pseudo-labeled data. For fairness, we also used the same 846 questions as the GeoTrust Data in our experiments.
We trained four models using the pseudo-labeled data and GeoTrust data with varying proportions (100%, 50%, 30%, and 10%). The trained models were subsequently tested on the GeoTrust-test set. As shown in table˜V, the model trained with 100% GeoTrust data slightly outperformed the model trained with pseudo-labeled data. This indicates that, given the same data volume, the synthetic data generated by TrustGeoGen achieves comparable or even superior results compared to pseudo-labeled data produced by SOTA models. It is worth noting that the SOTA model did not correctly answer all TrustGeoGen-generated questions, whereas TrustGeoGen can generate large volumes of data accompanied by accurate solution processes, providing a significant advantage for model training. Furthermore, the experimental results demonstrate that the model’s performance consistently improves with increasing training data size, further confirming the effectiveness of our generated data.
TABLE V: Performance comparison using different data augmentation strategies. "Pseudo-labels" and various percentages of "Our data" are compared against the "Pretraining" baseline. The $\Delta$ column shows the improvement over the baseline. Improvements are marked with an upward arrow ($\uparrow$).
Training data LLaVA-1.5-7B LLaVA-1.5-13B Qwen2-VL-2B Qwen2-VL-7B
Accuracy $\Delta$ Accuracy $\Delta$ Accuracy $\Delta$ Accuracy $\Delta$
Pretraining $1.67\%(4)$ - $2.08\%(5)$ - $3.33\%(8)$ - $4.58\%(11)$ -
[…]
GeoTrust data $10\%$ $5.83\%(14)$ $4.17\%\uparrow$ $7.50\%(18)$ $5.42\%\uparrow$ $3.75\%(9)$ $0.42\%\uparrow$ $11.25\%(27)$ $6.67\%\uparrow$
GeoTrust data $30\%$ $9.17\%(22)$ $7.50\%\uparrow$ $10.83\%(26)$ $8.75\%\uparrow$ $10.42\%(25)$ $7.08\%\uparrow$ $12.08\%(29)$ $7.50\%\uparrow$
GeoTrust data $50\%$ $9.58\%(23)$ $7.92\%\uparrow$ $11.25\%(27)$ $9.17\%\uparrow$ $10.00\%(24)$ $6.67\%\uparrow$ $12.50\%(30)$ $7.92\%\uparrow$
GeoTrust data $100\%$ $2.92\%(31) 11$$1.25\%\uparrow$ $14.17\%(34)$ $12.08\%\uparrow$ $12.50\%(30)$ $9.17\%\uparrow$ $14.17\%(34)$ $9.58\%\uparrow$
5.4.3 Advantages of by Thinking Template
We investigated the impact of our proposed thinking template across four model architectures. Our primary training dataset, GeoTrust-train, was constructed by augmenting a baseline dataset of 2,158 instances with 184 instances that incorporate the thinking template, resulting in a total of 2,342 instances. To isolate the contribution of the Thinking Template, we created three training setups:
  • GeoTrust-train (w/o TT): This version consists solely of the 2,158 baseline instances, with all Thinking Template data removed.
  • GeoTrust-train (Re TT): To maintain the dataset size for a fair comparison, this version replaces the 184 Thinking Template instances with 184 additional baseline instances, also totaling 2,342 instances.
  • GeoTrust-train: This is the full dataset including both the 2,158 baseline instances and the 184 Thinking Template instances.
Subsequently, each of the three resulting models was evaluated on the GeoTrust-test to assess their performance.
Fig. 7 illustrates the experimental results on the Thinking Template dataset. The Thinking Template demonstrates enhanced performance across all four models. Interestingly, for the LLaVA-1.5 series, the model trained with our Thinking Template (Re TT) consistently outperformed the version trained without it (w/o TT). In contrast, for the Qwen2-VL series, the model trained without the Thinking Template (w/o TT) achieved superior results.
Figure 7: Ablation Study of Thinking Template on GeoTrust-test
@@ 5.5 OOD Generalization
We evaluated the efficacy of TrustGeoGen in enhancing GPS performance by conducting extensive experiments across three diverse out-of-domain (OOD) benchmarks: GeoQA, Geometry3k, and OlympiadBench. These datasets were strategically chosen to cover a wide spectrum of challenges. GeoQA, sourced from real-world scenarios, offers diverse and unstructured problems accompanied by rich natural language rationales. Geometry3k provides formally structured problems generated via a formal engine, but without natural language explanations. To assess performance on highly complex tasks, we included OlympiadBench, which contains challenging competition-level problems. Together, this suite of benchmarks enables a robust and comprehensive evaluation of our approach.
5.5.1 GeoQA
The GeoQA dataset originates from real-world sources and comprises a high-quality, diverse set of problems presented in natural language along with their corresponding solutions. The dataset provides a training partition of 3,499 problems and a test partition of 754. In our experiments, we fine-tuned four distinct models via Supervised Fine-Tuning (SFT) utilizing three separate training datasets: the GeoQA training data, the GeoTrust-train data, and a combined dataset. Subsequently, we assessed the models’ efficacy on the GeoQA test set.
The results, presented in table˜VI, indicate that for the LLaVA-1.5 model family, all three training approaches yielded performance enhancements. It is particularly noteworthy that GeoTrust-train, despite being from a different domain, still enabled the models to achieve superior results on GeoQA. Moreover, a synergistic effect was observed when the two datasets were merged, with GeoTrust-train offering further incremental gains.
In contrast, for the Qwen2-VL model family, fine-tuning exclusively on the GeoTrust-train data led to performance degradation. A plausible explanation is that the Qwen2-VL series, being more recent, already possesses strong geometric reasoning abilities from extensive pre-training. This strong prior knowledge explains its superior initial performance, but also makes it susceptible to performance loss when adapted to a dataset with a distinct data distribution like GeoTrust-train. Despite this, using GeoTrust-train as a supplementary dataset for GeoQA still enhanced the models’ performance.
TABLE VI: Performance comparison on GeoQA dataset using different training data mixture. The $\Delta$ column indicates the improvement over the baseline. An up-arrow ($\uparrow$) indicates an improvement, and a down-arrow ($\downarrow$) indicates a decline.
Model Training Data Accuracy $\Delta$
LLaVA-1.5-7B - 13.66% (103) -
GeoQA 24.80% (187) $11.14\%\,\uparrow$
GeoTrust-train 13.79% (104) $0.13\%\,\uparrow$
GeoQA+GeoTrust-train 26.92% (203) $13.26\%\,\uparrow$
LLaVA-1.5-13B - 16.31% (123) -
GeoQA 25.46% (192) $9.15\%\,\uparrow$
GeoTrust-train 16.71% (126) $4.00\%\,\uparrow$
GeoQA+GeoTrust-train 30.11% (227) $13.80\%\,\uparrow$
Qwen2-VL-2B - 20.56% (155) -
GeoQA 25.46% (192) $4.90\%\,\uparrow$
GeoTrust-train 14.99% (113) $5.57\%\,\downarrow$
GeoQA+GeoTrust-train 27.72% (209) $7.16\%\,\uparrow$
Qwen2-VL-7B - 37.00% (279) -
GeoQA 39.66% (299) $2.66\%\,\uparrow$
GeoTrust-train 28.38% (214) $8.62\%\,\downarrow$
GeoQA+GeoTrust-train 43.50% (328) $6.50\%\,\uparrow$
GeoQA was evaluated in a direct question-answering (QA) format, rather than a multiple-choice setting. The same metrics as described in section˜5.1 were used.
5.5.2 Geometry3K
Geometry3K, a dataset constructed via a formal engine for solving geometry problems with formal language, comprises 2,101 training and 601 test examples. We performed SFT experiments on this data using different mixtures, and the results are summarized in table˜VII. For the LLaVA-1.5-7B, LLaVA-1.5-13B, and Qwen2-VL-2B models, every data mixture enhanced performance. Conversely, on Qwen2-VL-7B, training solely on Geometry3K led to a decline in performance. This performance drop persisted even when the training data was supplemented with GeoTrust-train.
Qwen2-VL-7B’s strong pre-trained performance on Geometry3K is characteristic of a well-saturated large multimodal model. This suggests that for such models, fine-tuning on data with a similar distribution to the pre-training set can paradoxically lead to performance degradation. Conversely, the smaller Qwen2-VL-2B model, whose pre-training was likely less saturated, still demonstrated further gains when fine-tuned on this data.
TABLE VII: Performance comparison on Geometry3k using different training data mixture.
Model Training Data Accuracy $\Delta$
LLaVA-1.5-7B - 1.16% (7) -
Geometry3K 6.66% (40) $5.50\%\,\uparrow$
GeoTrust-train 2.83% (17) $1.67\%\,\uparrow$
Geometry3K+GeoTrust-train 7.99% (48) $6.83\%\,\uparrow$
LLaVA-1.5-13B - 2.33% (14) -
Geometry3K 7.32% (44) $4.99\%\,\uparrow$
GeoTrust-train 3.33% (20) $1.00\%\,\uparrow$
Geometry3K+GeoTrust-train 8.65% (52) $6.32\%\,\uparrow$
Qwen2-VL-2B - 8.32% (50) -
Geometry3K 10.15% (61) $1.83\%\,\uparrow$
GeoTrust-train 8.65% (52) $0.33\%\,\uparrow$
Geometry3K+GeoTrust-train 12.15% (73) $3.83\%\,\uparrow$
Qwen2-VL-7B - 17.97% (108) -
Geometry3K 15.80% (95) $2.17\%\,\downarrow$
GeoTrust-train 19.13% (115) $1.16\%\,\uparrow$
Geometry3K+GeoTrust-train 16.14% (97) $1.83\%\,\downarrow$
Geometry3K was evaluated in a direct QA format, rather than a multiple-choice setting. The same metrics as described in section˜5.1 were used.
5.5.3 OlympiadBench-geo
OlympiadBench serves as an Olympiad-level benchmark comprising authentic competition problems to assess the multimodal reasoning capabilities of foundation models. For our evaluation, we designated a subset of 112 geometry problems from OlympiadBench as the test set OlympiadBench-geo to verify the efficacy of our proposed data. In the absence of a provided training set from OlympiadBench, we formulated several data compositions for our experiments by utilizing the training data from GeoQA.
The experimental results, presented in  table˜VIII, reveal a clear trend. Fine-tuning the four models exclusively on GeoTrust-train resulted in universal performance enhancements. However, for the Qwen2-VL-7B model, reliance on GeoQA data alone caused a decline in performance—a phenomenon likely associated with data distribution issues with its pre-training data, as discussed in previous sections. Crucially, combining the GeoQA and GeoTrust-train datasets produced marked performance improvements across all models, demonstrating their enhanced problem-solving abilities on this demanding benchmark.
TABLE VIII: Performance comparison on OlympiadBench-geo using different training data mixture.
Model Training Data Accuracy $\Delta$
LLaVA-1.5-7B - 0.00% (0) -
GeoQA 3.57% (4) $3.57\%\,\uparrow$
GeoTrust-train 1.79% (2) $1.79\%\,\uparrow$
GeoQA+GeoTrust-train 3.57% (4) $3.57\%\,\uparrow$
LLaVA-1.5-13B - 1.79% (2) -
GeoQA 6.25% (7) $4.46\%\,\uparrow$
GeoTrust-train 2.68% (3) $0.89\%\,\uparrow$
GeoQA+GeoTrust-train 6.25% (7) $4.46\%\,\uparrow$
Qwen2-VL-2B - 2.68% (3) -
GeoQA 4.46% (5) $1.78\%\,\uparrow$
GeoTrust-train 4.46% (5) $1.78\%\,\uparrow$
GeoQA+GeoTrust-train 6.25% (7) $3.57\%\,\uparrow$
Qwen2-VL-7B - 7.14% (8) -
GeoQA 5.36% (6) $1.78\%\,\downarrow$
GeoTrust-train 9.82% (11) $2.68\%\,\uparrow$
GeoQA+GeoTrust-train 8.93% (10) $1.79\%\,\uparrow$
5.5.4 Ablation Study of Thinking Template
As established in the preceding sections, our GeoTrust-train dataset is composed of two primary components: base data and thinking template (TT) data. To validate the effectiveness of the thinking template data, we conduct a series of ablation studies on the GeoQA and Geometry3K datasets. Following the same methodology outlined in section˜5.4.3, we partition the GeoTrust-train data into three distinct configurations for our experiments: without thinking templates (w/o TT), replace thinking templates (Re TT), and full data. Each configuration is then combined with the training sets of GeoQA and Geometry3K, respectively, to fine-tune the models. The models are subsequently evaluated on the corresponding test sets.
The experimental results on GeoQA, illustrated in fig.˜8, reveal that the inclusion of thinking templates consistently enhances performance across all models. For LLaVA-1.5-7B, however, the addition of more data, including thinking templates, yields only marginal improvements. This suggests that the model’s performance may have reached a saturation point with the base data alone. In contrast, LLaVA-1.5-13B demonstrates greater sensitivity to data volume; the performance gain from increasing the dataset size is more pronounced than that from introducing thinking templates. Conversely, the Qwen2-VL series benefits more significantly from the addition of thinking templates. This finding aligns with our earlier analysis that since these models are already well-pretrained, they are more responsive to novel data types than to a mere increase in data quantity.
The results on Geometry3K, as depicted in fig.˜9, present a different pattern. Notably, Qwen2-VL-7B achieves the highest score out-of-the-box (i.e., with its original pre-training), a result we attribute to its pre-training data composition. Nevertheless, when fine-tuned on our mixed datasets, the inclusion of thinking templates still provides a clear performance boost. Overall, the thinking templates demonstrate a more pronounced impact on the performance of the LLaVA-1.5 series models in these experiments.
Figure 8: Ablation Study of Thinking Template on GeoQA
5.5.5 Summary of Experiments Results
Our experiments OOD datasets and benchmarks confirm the generalizability of TrustGeoGen, demonstrating its ability to deliver consistent performance enhancements for models across diverse geometric datasets. A key component, "connection thinking", guarantees the logical coherence and variety of the synthesized data, to the extent that models trained solely on our GeoTrust-train dataset achieve notable improvements. The evaluated models display intriguing and divergent behaviors. For instance, the LLaVA-1.5 family—representing an earlier generation of models trained with limited geometric data and less sophisticated methods—shows considerable potential for growth. Our synthetic data consistently and significantly elevates its accuracy. Conversely, the Qwen2-VL series, which leverages state-of-the-art training paradigms and extensive geometric data, is already well-pretrained for these tasks. Nevertheless, enriching its training regimen with GeoTrust-train still yields further performance gains.
Furthermore, our novel "thinking templates" introduce structured, multi-faceted problem-solving paradigms previously unseen by these models during pre-training. This explains why even the robustly pre-trained Qwen2-VL models derive clear and objective benefits from the templates’ novel reasoning structures.
In summary, TrustGeoGen-synthesized data has demonstrated robust efficacy across a range of datasets and model architectures. Both of its core mechanisms, "connection thinking" and "thinking templates", have been proven effective. These results offer valuable insights and a compelling direction for future research on leveraging formal engines for automated data curation in training advanced models.